Factorise the expressions and divide them as directed. 
 (5p^2 –25p+20)÷(p–1)

Grade:12

1 Answers

Harshit Singh
2 years ago
Dear Student

Step 1: Take 5 common from the equation, 5p^2–25p+20, we get
5p^2–25p+20 = 5(p^2–5p+4)
Step 2: Factorize p^2–5p+4
p^2–5p+4 = p^2–p-4p+4 = (p–1)(p–4)
Step 3: Solve original equation (5p^2–25p+20)÷(p–1)
= 5(p–1)(p–4)/(p-1)
= 5(p–4)

Thanks